Actions from the Operational meetings

When raised Who What Comments Deadline
17/06/2020 Iain Make proposed changes to ILCC GPU server PSU configuration    
25/03/2020 Alison Support staff accessing user's own machines to debug problems. Make sure suitable speel is available to support staff  
27/02/2019 Alison/Jennifer Produce "get started" guide for new O365 users Mostly pointers to other documentation By time next group of users moves across
01/07/20 ALL Review "nograce" entitlements    
01/07/20 Alastair Chase whether Fortinet VPN bug has been fixed    
01/07/20 Craig Publicise advice for OSX AFS users to update client    
01/07/20 Adam Check new Uni VPN issues    

Blog entries

The blog is here.

When raised Who What Comments

Deferred actions

When raised Who What Comments Bring back
13/01/2016 All Units Remove unused kit from racks in IF, AT, KB   August
08/10/14 Services Trawl for pages restricted by ip address rather than via cosign   September (recurrent)
10/05/2016 George / Graham IPv6 for student labs Graham now implementing... September
13/01/2016 All Units Review rack use in IF, AT, KB   October (recurrent)
26/06/19 Alastair Ask IS how they manage server room incidents The University's out-of-hours policy is here Alastair is going to track down a name in IS. November
26/11/14 ALL units Review nagios use The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states. Refer to the Nagios Tactical Overview and Comments pages. November (recurrent)
11/01/12 ALL units Minimise rootmail   Dec (recurrent)
27/01/2016 All Review pages in computing.help Link to sorted-by-review-date list. January (recurrent)
13/01/2016 Tim/George On-line exams and EdLAN issues: schedule test Do after IPv6-for-on-line-exams; new EdLAN pending February 2021
26/11/14 ALL units Review nagios use The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states. Refer to the Nagios Tactical Overview and Comments pages. May (recurrent)
11/01/12 ALL units Minimise rootmail   June (recurrent)
08/10/2014 All Units Review pages in computing.help computing.help pages sorted by last review date July (recurrent)
03/11/15 ALL Review "nograce" entitlements   July (recurrent)

Completed actions

When raised WhoSorted ascending What Comments
12/06/13 RaT-unit Discuss security of exam scripts for external examiners with Neil M Not an issue - see 12/08/14
25/01/12 RaT-unit Try ColTeX on SL6 again Patch distributed
24/01/18 Alastair To investigate using Teams (rather than Slack) for alternative chat current status
27/02/2019 Alastair Raise issue of School office using Google calendar with Martin Project created to cover work
Structural Review Alastair Consider devolved budgets for units  
24/01/18 Alastair Discuss AT furniture proposals with Martin See specific suggestions in Section 4 of the Inf Unit Report for 8.8.2018
25/10/17 Alastair Propose to Stuart Anderson that students-in-grace should not have access to lab machines Stuart has agreed
Structural Review Alastair Consider devolved budgets for units Budget of 500/Unit agreed
23/08/17 Alastair Add "migrate to https" to Computing Plan  
22/02/17 Alastair Follow up "functional accounts" with Angus / Abdul Report scheduled for 26/4/17 CCPAG meeting. Toby will coordinate our eventual response.
10/02/2016 Alastair Security week Each group met to discuss priorities - See CEG report Alastair to write up notes from CEG meeting on 4th May. (being tracked at CEG)
27/04/2016 Alastair Printing, VAT - discuss with Martin  
27/04/2016 Alastair Attendance monitoring - Feed back concerns  
09/12/15 Alastair Discuss priorities and panic buttons with HoS Exams have priority
11/11/15 Alastair Arrange groups for and set date for security review Now next year
2015-09-23 Alastair College encryption policy  
2015-07-22 Alastair Chase up Campus agreement  
TDM 21/05/14 Alastair Investigate Mobile AV options further Transfer to techs though RT
25/02/15 Alastair Check status of SL7 partition layout  
12/02/14 Alastair Service Catalogue Take topic to CCPAG. Delayed until next CCPAG meeting  
09/07/14 Alastair Schedule/advertise "computing structure discussion" discussion Alastair to write up meeting
08/10/14 Alastair create wiki page with bullet points for CO structure meeting Done
08/10/14 Alastair Change title of holidays page subsequently done - wiki page created
12/02/14 Alastair Service Catalogue Review/remove fields in service catalogue Now being tracked at CEG
28/05/14 Alastair Clarify office options with Neil  
10/07/13 Alastair switchtoinfdb frequency? Frequency increased. Monitored and all seems well.
09/10/13 Alastair Take "security bollards" issues to Building Committee Security have agreed. Dave H has a key. Sheila is aware.
12/06/13 Alastair Take "power-down evacuation?" to Building Committee No action to be taken
28/11/12 Alastair Discuss wake/labs with DoT and Buildings Committee. Done. Spawned two new actions, RAT to update computing.help to to emphasis that students needing significant processing power for their projects should preferably use ECDF and MPU to consider additional wake features such as a command to wake individual lab machines
13/03/13 Alastair Investigate consequences of delegating some headers to other Schools MPU will create project. Will use git.
27/03/13 Alastair press for more say in when the building UPSes are turned off Done - to be turned back on 19/4/13
13/03/13 Alastair Investigate 32-bit Windows 7 laptops Most manufacturers supply 64=-bit Windows as a no-cost option.
13/02/13 Alastair Investigate whether we set up a secure location for examples of secondary authentication Now at SecondaryAuthentication
28/11/12 Alastair Discuss wireless with HoS.  
27/06/12 Alastair Ask IS for Office 365 IMAP instructions Alison circulated responses. Alastair to take to CCPAG
25/04/12 Alastair Take PGP/GPG discussion to CCPAG CCPAG Meeting 24/10
27/06/12 Alastair Discuss EE's public key infrastructure with Colin Higgs Now on to-do list
28/03/12 Alastair Set up USENIX membership Ordered
08/08/12 Alastair Check status and consquences of KB shutdown on 21st August  
08/08/12 Alastair Consider how to promote informal technical discussions amongst computing staff  
27/06/12 Alastair Raise issue of exam data policy with Liz Wait until college raises issue
27/06/12 Alastair Discuss how transportation of kit to KB should work with Dave H  
28/03/12 Alastair Email CCPAG about the possibility of cooperation between Schools in the event of disaster  
27/07/11 Alastair "BIOS settings" checklist SelectPCBiosSettings
12/10/11 Alastair Publicise auto-swap-size algorithm see here
12/10/11 Alastair Publicise auto-swap-size algorithm on lcfg wiki
12/10/11 Alastair Link T3 documents  
27/07/11 Alastair Raise BIOS settings for new machines at lcfg deployers meeting  
28/04/10 Alastair circulate grade 8 competency framework  
24/02/10 Alastair follow up on 24/02/10 "home directories" discussion  
11/02/10 Alastair Ask Dave Robertson and Gordon about Building Committee minutes Apparently not published (yet??)
13/01/10 Alastair Text messages to be sent to willing COs and policy on subsequent action to be decided. Take to CEG
11/11/09 Alastair Raise issue of forwarding homepages with Perdita  
24/06/09 Alastair Talk to Perdita and Dave Robertson about further steps we can take to promote social communication within the school  
25/02/09 Alastair Move network tester to 2.09 Done; please remember to sign it out when you use it
11/02/09 Alastair Investigate whether access to the installroot should be limited to HOST_MANAGED In release mechanism
22/11/17 Alastair / George Review and update the compromised machine instructions moved to CEG
12/06/13 Alastair/Chris College's data security pages and computing.help Our and College's data security advice pages are in accord.
11/04/12 Alastair/George Write letter about the continuing cooling issues in LV room. not required as issue resolved
30/05/12 Alastair/Ian IPMI for new server models  
14/11/12 Alastair/RAT Chase up Office365 IMAP with Dave B. RAT has tested and sent details to students.
28/08/19 Alison Speak to Iain about GPU desktops  
23/04/18 Alison IS unchaperoned server room access? Alison to document CEG decision with regards to this - see here
10/01/18 Alison Raise issue of Bayes support at CEG Bayes people are to be treated as Informatics staff for the present
23/11/16 Alison Ask Mohammed to mention mac-user mailing list when new kit goes out Done
2015-08-26 Alison Flesh out Campus documentation See Windows section
29/01/14 Alison www.anc/dtc Met Pim. www.anc/dtc will move to new Wordpress site on Mon 16th Nov
2015-09-09 Alison Email returning students re SL7  
25/02/15 Alison Organise CO talk at jamboree  
09/04/14 Alison put pandemic mail list information on Pandemic pages  
12/02/26 Alison Windows XP - end of life Take to CEG to discuss policy
29/01/14 Alison Innovative learning week See Informatics Innovative Learning week and US Unit report
12/02/14 Alison staffmail vs exchange Glitch last August was not procedural i.e. not due to human error
13/11/13 Alison Review NotifyUserDisruption page Done
26/09/12 Alison Email ITPF re UniDesk wiki page to incorporate Angus's suggestions and any other useful contacts
11/09/13 Alison RT best practice Document updated, Alison to mail out - RTBestPractice
24/07/13 Alison Review old *-team lists  
24/04/13 Alison Talk to Neil McG about ISSRT ticket retention time Will apply retention times on a queue by queue basis. Alison to email details to George/Iain - being tracked at CEG
26/09/12 Alison Issues regarding IS's handling of UniDesk tickets Aggregate 'in-house' fixes - see UniDeskTips
08/05/13 Alison Talk to techs about their RT queue  
27/03/13 Alison Change MPU RT mail settings to allow all Unit members to see correspondence Done - tested after 10/4/13 meeting
16/05/12 Alison Update wiki page on procedures on safe disposal of disks Procedure Document
12/10/11 Alison Add student password portal concerns to CEG action See also the wiki page
08/02/12 Alison Convert document about updating users into a checklisk NotifyUsersNotes
12/10/11 Alison Add student password portal concerns to CEG action Done and discussed
11/01/12 Alison Create a wiki page to note password-entry locations PasswordEntry
23/11/11 Alison Write procedure for contacting users about unscheduled service disruptions Done - see here
26/10/11 Alison Discuss visitors' (non)-sponsors with HR  
12/10/11 Alison Discuss pcounter credit refunds with ISS  
31/08/11 Alison Write up relationship with IS helpdesk To be done as part of DevProj:177
14/09/11 Alison check that Stuart has sent e-mail to students explaining new printing charging  
26/01/11 Alison Create a development project for improving the roles mechanism DevProj:197
10/11/10 Alison Add notes on syntax to location tag files Thanks to Ian D
10/11/10 Alison Stephen suggested that date in 'when raised' field should link to the minutes  
28/04/10 Alison take concerns over new web site to health and safety committee  
28/10/09 Alison Create wiki page to record machine problems requiring reboot Linked from us-unit wiki page
14/10/09 Alison Check both invquery and minv provide accurate info. They don't but will be fixed as part of inventory project
24/02/10 Alison (documentation project) follow up on 24/02/10 AFS discussion Tracked as part of documentation project
14/08/13 Alison/Craig/Stephen DES documentation - blogs/web page instructions by OS  
09/06/10 Alison/George Investigate AT server inventory issues script now amended
12/11/14 alisond set up meeting about visitor accounts Tuesday 2nd at 14:00 in IF-4.02
11/12/13 alisond Publish checklist for firewall holes Circulated to CEG and comments incorporated. Now circulate to COs. Done
11/12/13 alisond Circulate slides on account creation process Linked in General Information section - UserSupportUnit
14/01/09 alisond Investigate severe outage user communication options Done
09/01/19 ALL Pandemic Planning We Name The Guilty!!!!!! - now completed! (22/01/20)
29/01/14 ALL Consider festival of creative learning week removed from Action list until further notice (22/01/20)
03/11/15 ALL Review "nograce" entitlements Done - Services, US, Inf and MPU
27/03/19 ALL Consider which further entitlements should be removed from the staff role  
27/02/2019 All Mail Stephen with comments/suggestions about recent data incident Meeting arranged for end March
29/01/14 ALL Consider festival of creative learning week  
27/01/2016 All Review pages in computing.help Marked as completed before next reoccurrence comes along in July
25/04/18 All Machines in rack 2 managers of these machine liaise with each other to spread the load across both banks.
14/02/18 All Try out the replacement NX service see computing.help  
11/04/18 All Move/delete data from nix:/disk/huge Done - thanks to Neil
09/08/17 All All SL7.2 servers to be upgraded to SL7.3  
29/01/14 ALL Consider festival of creative learning week  
28/06/17 All remove sl6 profiles that are no longer required by end of August. Stephen will send out updated list
23/08/17 ALL Review InformaticsFunctionalAccounts page, and update as required  
08/02/17 ALL Add new and past console funnies to the wiki page CrazyConsoleAKABrownEffect
11/11/15 ALL Read Toby's roles-management problem statement TheProblemWithRoles ServicesUnitNograceEntitlements
12/10/2106 ALL Ensure SL6.8 upgrades are done Including kernel reboots as necessary - see SL66Profiles
29/01/14 ALL Consider innovative learning week There seems to be some doubt about when this is (and what it's called now)
12/10/2106 ALL doodle over Christmas lunch Details Deposit to Neil please
14/09/2016 All Consider importance of static IPv6 addresses  
27/04/2016 ALL Clear out old SL6 profiles Stephen has nagged about sl6 profiles SL6OfficeMachines. Support will investigate non-CO machines
09/12/15 All Reduce noise in lcfg status reports  
2015-08-12 ALL Comment on advice on Internet safety now published
27/04/2016 ALL Server specification Re-read and comment on the document
13/01/2016 ALL Think about KVM servers and VMs with end users Here's a discussion document: On the segregation of KVM guests All to read for next Operational Meeting
13/01/2016 All Tidy inappropriately-develop machines  
27/01/2016 ALL Check/label kit in B.03 Anything not claimed by end-February will be assumed to be no longer required.
2015-08-12 All Start using new VPN configurations See blog. Old configurations turned off on 2nd November
25/03/15 All Help out with 5th May exams if possible Roger and Neil offered to help
Structural Review ALL Check and update telephone contact list Done
25/07/09 ALL Send partition layout thoughts to Stephen  
09/04/14 All replace SSL certificates and restart affected services  
12/03/14 ALL Consider ramifications of JCMB management subnet renumbering  
25/09/13 All Think about how to make service catalogue more useful before meeting on 12th February All to add usage cases to ServiceCatalogueBrainStorm - discussed 12/02/14 and further actions added to list
25/09/13 All Move to using rmirror-client.h header ServicesUnitMirrorService and Neil's script - just venus remaining and it's due to be turned off. Suggest we say this is done.
11/09/13 ALL Comments ASAP on JCMB aircon proposals Received - thanks
12/12/12 All Note down occasions on which secondary authentication should be required. See SecondaryAuthentication - MPU to create project
28/11/12 ALL Reboot ASAP for kernel updates. only catzilla left - completed Fri 19th Apr
12/09/12 ALL Remove unneeded machines from AT server room venom now removed from rack and placed in B.03
12/09/12 ALL Move managed machines from SMSR to main server room Ian asked that this be dropped from here - 10/4/13
27/02/13 ALL upgrade remaining SL6.2 machines ... or take out of service
24/10/12 ALL Identify and label all 'spare' machines and remove 1850s. Identify bids for old kit. There's a ParkedKit page. Labels have been added to spare machines in Forum. Units have bid for old kit.
12/12/12 All Note down occasions on which secondary authentication is required. See discussion items
14/03/12 All Move some reusable retired kit to AT server room This referred to the idea of parking them on some Dexion shelving
30/05/12 ALL Restart VMs on central as required  
23/11/11 All Identify which servers should have local home directories MPU have created a web page listing the machines
09/11/11 All All units/service managers to consider which of their services are to are to be available to guest wireless users, and then to make the necessary arrangements. Inf Unit can advise. Superseded: all Informatics services should now be accessible by 'guest' wireless users. See Infrastructure Report, 23.11.20121, point 4 and Infrastructure Report, 14.12.2011, point 3
26/10/11 ALL Feed back to inf-unit re JCMB EdLAN-link work ASAP!
28/09/11 All Restart VMs Enough done for now
31/08/11 ALL Roles changes Went live 1st September!
17/08/11 ALL Think about IS helpdesk  
18/05/11 All Remove redundant LCFG reminders  
22/06/11 All see which pages, if any, are still relevant in the old groups documentation tree  
24/11/10 All Each Unit to trawl through bugs in bugzilla.inf not applicable for MPU as used only for weekly release
12/01/11 All feedback comments to George on Compromised DICE machine policy  
24/11/10 All Fix cabling in rack 3 , Forum server room Done
24/11/10 All Review and update pandemic documentation Done apart from Database which is still in a state of flux
24/11/10 All Fix cabling in rack 3 , Forum server room  
13/10/10 ALL Review and act on Neil's non-CO-login list For amd removal - US done, will be in stable on Thursday 2nd December
13/10/10 ALL Review and act on Stephen's "devel" bucket lists  
11/08/10 All SL5.5 servers to be rebooted by end of August Still waiting on one RAT machine
14/07/10 All Answer Ian's questions in Infrastructure Report  
23/06/10 ALL Consider Craig's suggestion of doing the KB EVO firmware upgrade during the Forum power-down Agreed that this is a good plan
09/06/10 All Units to consider their future server needs  
12/05/10 All Forward suggestions for additions to self managed server room web page  
14/04/10 All reboot servers requiring SL5.4 updates  
13/01/10 All Decide on an appropriate procedure for shutting down self-managed servers Discussed at 10/03/10 meeting
13/01/10 All Define the criticality of machines Now waiting for mp-unit work
28/10/09 All Start planning move from FH Complete!
13/01/10 All Ensure that all machines in the server room have their corresponding entries in the fpdu/sxx.outlets maps labelled correctly  
13/01/10 All Investigate best use of temperature monitoring. Part of cc project
13/01/10 All Test out the new query tools as per RAT report for 13/01 completed
11/11/09 All Complete pandemic actions to be tracked at CEG
28/10/09 All Move kit from BP  
09/09/09 ALL Reboot machines as soon as possible for latest kernel  
23/09/09 All Consider Neil's discussion point about exporting group data  
12/08/09 All Reboot servers to pick up SL5.3 by end of August Overtaken by latest kernel
24/06/09 All Consider whether it will be practicable for us to turn off AMD on servers once the AFS based RPM service is available Merge with other AMD action
22/07/09 All Check cron logs. Checksum problem with hostname has caused lcfg-cron configure failure in some instances No longer necessary
22/07/09 All Consider plans for moving kit from FH Wiki page available
08/04/09 All Move kit from BP and FH - aimimg towards 2 sites + off-site at JCMB merged with another action
08/10/14 All Units review pages in computing.help Everything prior to Jul 2018 - computing.help pages sorted by last review date
26/06/19 All Units Complete upgrade of remaining servers to SL7.6 Done - Services, US, MPU and Inf have 2 to go
26/06/19 All Units Complete upgrade of multi-user servers to SL7.6  
27/03/19 All Units Check which services depend on the Staff role and Linux group Services - Now all done
26/09/18 All Units Have servers upgraded to SL7.5 by end October  
13/01/2106 All Units Clear out unwanted machines from racks in IF, AT, KB  
10/01/18 All units Check whether web sites under their control are still using Allow/Deny configuration directives Switch was made, not much seems to have broken
10/01/18 All units Move all servers to SL7.4 Deadline amended to end of April 2018
28/03/18 All units Prepare reports for scheduled JCMB and Forum power-down reminder email sent out 28/03/18
13/09/17 ALL Units List "http:" sites Neil will create project to cover move of sites to https
11/01/12 ALL units Minimise rootmail Not sufficient to worry about
26/11/14 ALL units Review nagios use. The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states.
26/11/14 ALL units Review nagios use. (Recurrent action)
11/01/12 ALL units Minimise rootmail  
13/01/2106 All Units Review rack use in IF, AT, KB Old kit can be put in B.03 but please ensure serial number or name of machine is clearly visible. Services to arrange removal of old tape library
11/01/12 ALL units Minimise rootmail  
27/05/15 All units Review IF rack usage See ForumRackPopulation . See cos emails dated 8/7 about USB ID dongle information
2015-06-24 All units Review rack use in IF, AT, KB    
13/05/15 ALL units Review nagios use The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states. Refer https://nagios.inf.ed.ac.uk/nagios/cgi-bin/tac.cgi
27/08/14 All Units all servers to be running SL6.5 by end October Nearly there!
26/11/14 All Units Review nagios use. The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states.
11/06/14 ALL units Minimise rootmail  
24/07/13 ALL Units Procedures for penetration testing to be tracked at CEG
24/04/13 All units Move junked machines to B.03  
25/04/12 All Units Upgrade sl6.1 servers to sl6.2 Deadline end of May - expect deadline will be met. It was.
28/03/12 All Units/US Work through cookie list identifying cookies for which they are responsible  
26/09/12 ALL/Ian nagios oddities See Infrastructure Report  
2015-09-23 ALL/Neil/Alison Christmas lunch suggestions discussed 14/10/15 - Neil sent out doodle poll
26/06/13 ALL/US Review "to-be-kept" contents of B.03 US will ask Units about kit and if to be kept, will record reason. Alastair will have final say.
22/10/08 ascobie Approach School re contribution to Christmas lunch Done
22/10/08 ascobie Put holiday policy on CEG to-do list Done - on CEG wish-list
08/10/08 ascobie Investigate availability of central bulletin board service Done - the service has been suspended indefinitely
10/09/08 ascobie Find holidays policy page Done
10/09/08 ascobie Look at whether can add remote USB connector for SmartBoards onto lecterns being tracked at meeting room committee
13/08/14 AT dwellers Agree room occupation for level 8 Done
13/09/17 Carol Sweep of FH for junk A14 now clear of all Informatics stuff. A15 has had an uplift from CCL. Some of the remaining desktops will be going to charity and some need to be brought to the Forum
25/10/17 Carol Amend decommissioning procedures to add removal and storage of disk brackets. There is now a large purple tub in B.03 labelled "Disk caddies for servers"
25/10/17 Carol Investigate access rfe access requirements of technicians Technicians emailed. Gibert has responded saying that he needs access to the switch files. Garry responded saying he also needs this access plus uses rfe to look at lcfg profiles of machines.
10/10/2018 CEG Review initial 30-day account-use period 45-day period agreed and implemented
2015-08-26 CEG Consider reallocation of older kit  
Structural Review CEG Formalise and resource support for self-managed platforms Discussed and being tracked at CEG
23/11/11 CEG Record arguments for and against local home directories On the CEG wish list
Structural Review CEG? Organise another documentation week Pencilled in for November 2015
01/04/20 Chris Email out another reminder about remote working Done on 02/04/2020.
12/06/19 Chris Produce script to allow toohot to be easily used by self-managed machines Done amended action - mailed selfmanaged-sr with howto details
27/02/2019 Chris Add reference to Overleaf to computing.help computing.help.inf.ed.ac.uk/overleaf
27/01/2016 Chris Investigate Virtual DICE diffs Done, and linked to from here
2015-09-23 Chris lightdm WM choice menu image Updated (picture).
11/03/15 Chris Check the toohot component on Dell 730s Done
10/12/14 Chris Find home for dumpdep notes Where's my software?
12/06/13 Chris Run fs flushall before DICE sleep Sleep now runs flushall
12/06/13 Chris Link new "social media" guidelines into computing.help Now linked from the Guidelines page.
12/06/13 Chris Should lab 780s and 790s sleep or not? They will sleep w.e.f. the stable release of 17/7/13
27/02/13 Chris Merge remaining documentation items from ToDo list into RT queue ... or at least gather everything together into one location.
28/03/12 Chris Move Nagios code from individual hardware headers to server.h and small-server.h  
12/08/09 Chris Mail out details of proposed cron job start time changes Done
12/11/14 cms update Pandemic planning page to say that people should add themselves to appropriate mailing lists  
11/12/13 cms Update pandemic planning page with rationale for exercise  
31/08/11 Convener Take iDEA-lab-type web sites to CEG for guidance  
10/09/08 Convener Check/Redo Operational meeting room bookings Done until end of year
10/09/08 Convener Raise "Server Decomissioning Procedure" at CEG Done
10/09/08 Convener Raise "Disaster Recovery Procedures" at CEG Done
10/09/08 Convener Raise "server reboots for 5.2 upgrade" at CEG Done
25/03/2020 Craig Investigate more effective use of Teams for on-line meetings  
11/09/19 Craig Clear spurious bookings for AT-7.14 Done - we don't appear to have had any other bookings for this room.
12/06/19 Craig Take issue of server room monitoring back to CEG  
12/06/19 Craig Discuss with Alastair how to expedite load-shedding Alastair meeting with Jane on 13/6/19 to discuss
27/03/19 Craig Raise issue of more transparent recruitment at CEG  
27/02/2019 Craig Raise issue of ISS using Google calendar with Neil H See above
10/10/2018 Craig Check password portal wording  
23/04/18 Craig Take unmaintained personal web pages to Web Strategy Group  
28/09/2016 Craig Document Auristor client for self-managed Windows machines auristor
25/01/17 Craig Check whether Tibs 3 client works with TiBS 2 server and whether this would fix sierra issue It doesn't. There's a project to cover the TiBS3 upgrade
27/07/16 Craig Discuss MDP/MFD printer accounting with Martin  
23/03/2016 Craig investigate redirecting autofs debug output away from syslog.  
09/12/15 Craig On-line exams and pandemic list? Documentation here
2015-08-12 Craig Take IPMI issue to CEG MPU and Inf Unit will discuss
12/11/14 Craig review password portal rewording Done
2015-06-10 Craig Create a jabber project Done
10/12/14 Craig Take issue of copyright images on webpages to Web Strategy Group Sub-group to be set up by Steve Scott
TDM 21/05/14 Craig Further investigation of mobile printing To be submitted as a project. Done.
10/12/14 Craig Take issue of copyright images on webpages to Web Strategy Group  
26/11/14 Craig Mac mini and VirtualBox. documentation
08/10/14 Craig review pages in computing.help Procedures for use with the new Mac Mini - merged with other mac mini action
10/12/14 Craig Move actions from computing structure discussion discussion to operational meeting done
09/07/14 Craig Tape library dust filters? Doesn't have them
09/07/14 Craig Poll for TDM options Last Wednesday afternoon of month
11/06/14 Craig Document use of CO iPad Link UsingTheCoIpad somewhere
09/04/14 Craig test volume release script with x86_64 version of perl::AFS  
24/04/13 Craig Ask Sheila to check HP maintenance contracts As we thought, all servers have 4 years on-site warranty
27/03/13 Craig talk to Sheila about what our HP warranties provide Done - they provide on-site warranty
23/01/13 Craig Sort out management of websites hosted on dunkvm Kenny Bell is taking over management of websites but will need help. Craig will discuss with Colin Adams
12/09/12 Craig Raise MFD-printing evaluation issues upstream Craig has emailed Angus Rae
14/03/12 Craig Raise possibility of cooperation between Schools in the event of disaster at the next Scicos not-a-conference Raised at CCPAG. In view of other Schools lack of resources, we will rely on IS's virtual provision
28/03/12 Craig Trawl through user and group space for cookies Take to HoS; write report - tracked at CEG
14/12/11 Craig Edit blog article to include Inadvertent disk shutdown doc  
30/05/12 Craig Arrange meeting re amd maps meeting held
11/01/12 Craig Update docs to warn against some characters in passwords  
11/01/12 Craig Check how password portal handles metacharacters Inconsistently! Fixed.
26/10/11 Craig Amend password portal character class level  
14/09/11 Craig check printing refund process  
31/08/11 Craig AT-printing brain-dump and US training  
23/03/11 Craig Create web page with procedure for announcing interruptions to services See ServiceInterruptionProcedure
09/03/11 Craig Check out tapes in B.03  
26/01/11 Craig Bring backup retention proposals to CSG  
26/01/11 Craig Create a development project for migrating School mailing lists to the central server devproj:193
26/01/11 Craig Create a development project to look at archiving DevProj:194
10/11/10 Craig Create project to convert services-unit servers to apacheconf will be done by 6th December
22/09/10 Craig Clarify whether CSBE printing uses our printer wire It doesn't
12/08/09 Craig take the proposals for a new general purpose Informatics mailing list to Dave Done
11/03/09 Craig Investigate MFD costs Action adopted by CEG
13/09/17 Craig / Toby Discuss purge of old AFS admin IDs  
28/11/12 Craig/Graham maelcum upgrade. Graham doing it
17/08/11 Craig/Neil Follow up on printing issues  
12/03/14 Craig/Toby investigate progress of AFS patch at conference In Gerrit now, will be in 1.6.8
22/07/09 Craig/Toby Thrash out final solution for keytabs Done
11/12/13 gdmr Report recent compromise to Brian Gilmore Done. No response.
13/10/10 gdmr Prometheus -> CEG pandemic list  
10/09/08 gdmr Fix routing component to allow static route to be added via LCFG resources https://bugs.lcfg.org/show_bug.cgi?id=111
13/08/08 gdmr Potential gap in off-site provision Taken to CEG
13/08/08 gdmr Unblock IRC ports Done
27/06/18 George Create project to cover actions arising from server room notification investigation DevProj #558
24/04/13 George Fix the wire_*.h headers Only the Forum managed DICE wires remaining.
28/06/17 George SLAAC-style forward DNS for server wires All now done. A handful of non-server wires remain.
10/01/18 George Create no-slaac option for dns/inf6 rfe map See the inf-unit report
10/01/18 George Create project to address deprecation of legacy style rsyslogd rules #452
22/11/17 George Review rsyslog rules Easy part done. George to create project to cover hard part
13/12/17 George Find out if there's a spare corner of the KB College server room for the old tape library There is
22/02/17 George /netmask versions of subnets header files live/subnets_nm.h (see the inf-unit report)
26/10/16 George Core switch firmware See the inf-unit report
27/07/16 George IPv6 on wire M Done
09/12/15 George "external connectivity" computing.help page It's here. Suggestions for improvement welcome!
2015-10-28 George Add iptables to SL7 desktop project Done, with a mid-December deadline
11/09/13 George Infrastructure for commercialisation decant from AT Rolled into general AT decant
27/11/13 George Ask Dave about about sniffing data to extract URLs hosted on self-managed machines Dave has agreed. Note, though, that under the LBP regulations we have to publicise that we're going to do this before we do it. (We also have to publicise re the IDS project and the potential sFlow project.)
11/09/13 George CEG to discuss mod_waklog Discussed 24th September. MP-unit to take mod_waklog.
12/06/13 George Check HV restart with Mike R  
22/08/12 George Set up aliases for DHCP servers ifdhcp, atdhcp, kbdhcp
22/08/12 George Name and shame ancient AT kit  
28/03/12 George Investigate new rack for AT server room RT:57384
11/01/12 George Add more debugging info to lcfg-routing  
26/10/11 George Chase netflow data from Sam
01/07/09 George Document how the host keys of the ssh servers can be preserved during upgrades Subsumed by wallet project
12/09/12 George/Alastair Replacement of Forum switches  
14/01/15 George/Alison Set ports in all Forum meeting rooms to Conf110 apart from existing DICE ports which should be MAC-locked Done.
10/07/13 George/Alison Take "self-managed machines with firewall holes" to CEG See CEG minutes for August 12th
2015-08-26 George/ALL Set up meeting to discuss internal documentation FutureofInternalTechnicalDocumentation Action kept until week scheduled
30/05/12 George/Craig/Ian JCMB server room cleaning Scheduled for 18th June
25/07/09 George/Ian JCMB dust protection? All part of master plan
TDM 21/05/14 George/Toby/Neil Documentation for VPN on iOS/Android Android - draft
28/09/2016 Gordon MacOS Sierra Only remaining issue is lack of TiBS support
2015-09-09 Gordon Check SL7 printing to IS "charged" queues Done.
23/11/11 Graham Generated mailing lists and grace periods GracePeriodsMeeting2013 - project proposal incoming - now being tracked in Unit
09/04/14 Graham provide Craig with information on how to change mailing list memberships, roles etc in Theon to allow pandemic people to receive mail, nagios alerts etc Done
09/04/14 Graham address issue of accounts being created for applicants  
09/04/14 Graham come up with scanning tool for SSL issue  
12/03/14 Graham Provide Services Unit with patch for role rfe map delete script  
10/07/13 Graham Discuss snowmen with Toby and Neil  
08/05/13 Graham Deploy Crichton to develop machines  
27/03/13 Graham distribute notes from the recent UKUUG spring conference see here (Graham) and here (Chris)
14/11/12 Graham "post-package-search" script. Done
27/06/12 Graham File enhancement request about handling branches in source code  
28/03/12 Graham Investigate issue of user crontabs further done - but going to write blog
26/01/11 Graham Action automatic generation of the sys-announce mailing list  
10/06/09 Graham mail out with details of autoreboot mechanism Ongoing
13/03/13 Graham/mp-unit File enhancements to certain components to deal with USB/VirtualBox issues Tracked by Unit
10/01/18 Iain Investigate setting up disc zeroing station  
24/01/18 Iain review School Data Protection statement  
10/02/2016 Iain exFAT legality awaiting guidance on signing of NDA
23/04/14 Iain Beowulf and GPU-machine shutdown when cooling fails See Hadoop Care and Feeding.
12/06/13 Iain Revamp cluster shutdown scripts Scripts done. Now install them somewhere and write some docs. See HadoopCareAndFeeding
10/07/13 Iain Follow up on e-expenses registration IS responded to say that they intend implementing a random password generator. Using EASE is not a simple option.
08/05/13 Iain Email out details of changes to default RT quicksearch behaviour  
22/08/12 Iain Take SL6.3 decision to RAT for consideration  
14/12/11 Iain Organise meeting about DHCP enhancements Done
14/12/11 Iain Meet with Beetle project to discuss issues preventing upgrade of their machines to SL6  
14/09/11 Iain check version of webots on lab machines and send out a suitable e-mail  
28/04/10 Iain report back on progress of FC12 port redundant
28/04/10 Iain create bugzilla actions for FC12 port  
12/05/10 Iain make platspec header entries explicitly SL5  
13/11/2019 Iain, Support Improve documentation for GPUs in labs. Confirm access to machines and access to level 9 Chris is producing something
27/03/13 Iain/Alison look into making RT mail behaviour more consistent. By default all queues will notify Unit on ticket creation and when ticket moved to queue. If a Unit wants all members of the group to be watchers on the queue, contact Alison
11/04/12 Iain/Alison Demonstrate RT4 Will be rescheduled for some future slot
14/08/2019 Iain/MPU Investigate issues with cluster file server installation Iain to contact MPU
13/05/09 Iain/Toby/Alastair Take "certificates" to CCPAG. No longer needed
28/01/09 iainr move beowulf profiles to normal LCFG servers and decomission illustrious Done
13/08/08 iainr Collate responses re proposed University's Source Repository Done, being tracked at CEG
10/02/2016 Ian SMSR power down - Create an 'Emergency' page on computing.help See https://computing.help.inf.ed.ac.uk/smsr-emergencies and http://computing.help.inf.ed.ac.uk/smsr-ups-monitoring
25/09/19 Ian Talk to School procurement about investigating the power requirements of prospective purchases
13/03/19 Ian Investigate IPv6 and Nagios Done. See the Nagios and IPv6 item of the Inf Unit Report, 27.3.2019.
25/04/18 Ian write a report on what is happening in AT server room Done - see email to cos 3/5/18
24/01/18 Ian Investigate how to best test IPv6 firewall holes see minutes
27/07/16 Ian Document current serial console use Done: RS232SerialConsoleProvision2016Review
27/07/16 Ian Speak to IS re ESSIS replacement Done: no action required by us; some problems/issues (namely: limitations on allowed number of users; deficiencies in the scheduler) with the previous service have been fed back.
26/06/13 Ian tlsdir locations Now a X3 round-robin
26/06/13 Ian Mirroring and retention of console logs Now mirrored to loghosts
23/11/11 Ian Investigate why nagios did not report the fibre problem sooner Done - see email sent to COs on 24.11.2011
28/04/10 Ian look at splitting UPS mailing list Mailing list not split; instead, KB UPSes reconfigured to send email alerts as necessary every 12 hours, not every 2 minutes
14/04/10 Ian Clear shelf space in Forum server room  
28/10/09 Ian Look at power management for machines with IPMI v1.5 Documentation
10/02/2016 Inf NUT service for SMSR Now an Inf Unit project
22/10/14 Inf Compile notes on extracting firewall information See our "how the network filtering works" document and our 2014-11-12 report
13/11/13 Inf ask IS what actions they are proposing to take re signing ed and its subdomains See inf-unit report. Expected in F/Y 2015-16
14/03/12 Inf Record locations of RFE maps Linked at the end of the netmon pages
28/09/11 Inf Investigate rotation of Jabber chat logs Answer: currently done by a daily cron job (set up by live/jabber-server.h) on curlew which deletes any 'cos' chat room logs over 90 days old. The facility for chat room log rotation/culling appears not to be provided by the Jabber suite itself; the only configuration choice for any chatroom seems to be whether or not it is actually logged or not - if it is, the logs persist until external action to rotate or cull them is taken.
28/09/11 Inf Investigate non-CO Jabber usage Part of DevProj:221
28/09/11 Inf Produce proposal for update to Jabber service DevProj:221
28/09/11 Inf Produce document gathering subnet definitions together live/subnets.h
13/01/10 Inf Don't advertise private IP addresses outside Informatics is less trivial than first expected but progress made. Forthcoming upgrades to BIND may make things easier
22/06/11 Inf Add investigating weak password detecting PAM module to Devproj #168 - Password strength checks  
10/11/10 Inf Suggested that there should be a pinboard in server room Done, 11.3.2011
13/10/10 Inf Write a "what do do when things go wrong" document posted in server room
09/06/10 Inf Review what's currently connected to fibre channels and investigate load balancing See the Inf Unit report, 14.7.2010
11/09/19 Inf Unit Chase up problems with AT server room alarm keypad  
11/09/19 Inf Unit Provide Person At Work sign for AT server room  
24/07/19 Inf Unit IPMI password length bug RT#96635
27/02/2019 Inf Unit Send out general warning about AT server room power work  
12/03/14 Inf Unit Chase old self-managed machines decomissioning action RT#67343
22/09/10 Inf,US Units to record system criticality using Neil's macros All now complete
10/07/13 Inf-unit Alternatives to /admin Ian to bundle into DevProj#279
28/11/12 Inf-unit Try to track down bogus SSIDs.  
12/10/11 inf-unit Add "how to disable" etc comment to .ports files Done, 19.10.2011
06/04/11 inf-unit Check self-managed server room noise levels  
10/03/10 inf-unit Front-page for self-managed server room Initial skeleton
24/02/10 inf-unit document bonding issues see here
13/05/09 inf-unit Mail out to if-people re. wireless Done
11/03/09 inf-unit console server provision in self-managed server room Done
12/11/08 inf-unit Arrange tools for server rooms Done
10/09/08 inf-unit Coordinate forum server room rack rotation and floor grill replacements Done
23/07/08 inf-unit Wireless in the Forum Done; RT:37791
13/08/08 inf-unit Various Forum server room things Done
23/07/08 inf-unit Unify naming of switches with floorbox ports Done, Feb 2009
13/03/13 inf-unit / mp-unit Discuss dhcpd component changes ProposalsForChangesToLcfgDhcpComponent; Inf Unit and MPU to finalise proposal; take outcome to LCFG deployer's meeting
09/03/11 Inf/MPU Investigate NTP clocksource problem further Rebooting fenrir fixed problem
26/05/10 Infrastructure Unit Propose scheme for root passwords  
22/10/08 ktd Put actual number of tickets in RT resolution reports Done wef 12/11/08
13/09/17 MP Unit Reconsider "real" serial consoles for KVM hosts  
28/06/17 MP unit Post mortem on KVM server crash Results inconclusive
26/05/10 MP Unit Consider how best to use the new sysinfo.criticality resource, following discussion on 9th June see minutes from 28/07/2010 for update
11/09/13 mp-unit inf-level machine for package testing Done, see the inf layer
12/06/13 MP-unit Sundry "criticality" things from discussion  
25/01/12 MP-unit Remote access arrangements to COs' desktops DONE - dice/options/co-desktop.h
14/04/10 mp-unit CUPS not to be running on servers by default More than just servers?
10/03/10 mp-unit Document power-button handling options  
24/02/10 mp-unit add profile resources for machine criticality On mp-unit's list
11/03/09 mp-unit Look at om disowning credentials Done. Fix in release cycle
28/01/09 mp-unit Change DIY DICE root password Done; ask Alastair if you need to know it
10/09/08 mp-unit Make DBAN easier to use On MPU to-do list
26/10/11 mp-unit/All Add nagios hardware checks to some machines by default Done, but see follow-up action
22/01/2020 MPU Chase kernel upgrades  
25/03/2020 MPU Create header file for Dell PE740xd2 Will be in stable release next week.
01/04/20 MPU Update MOTD on XRDP servers  
24/10/18 MPU Add banners to XRDP hosts if possible
10/01/18 MPU Document procedure for returning failed discs to Dell Done
13/12/17 MPU Review procedures to be followed when a machine goes self-managed, particularly the deletion of the DICE host principal Wiki page created
26/07/17 MPU produce script to monitor package volume usage  
26/07/17 MPU Adjust quotas for package volumes  
25/01/17 MPU Produce script to fix "dracut" issues Stephen ran cron job
28/09/2016 MPU Set up script to mail out monthly firmware update reminders Now on MPU to-do list
28/09/2016 MPU Add investigation of SL7 multi-homed issue to to-do list RT#79713
08/06/2016 MPU Look into issue of KVM guest segregation and present recommendations Recommend a separate server for non-CO access.
27/04/2016 MPU Logging and reporting of unexpected USB events? included in Security week action
27/04/2016 MPU HP 800 G2 computing.help and/or BIOS pages Transferred to MPU list
2015-09-23 MPU Add encryption of tmp and swap to <develop> Done (in stable of 11/11/15)
2015-09-09 MPU DICE_STICK_WITH_SL65?  
2015-08-26 MPU Investigate usage of pkgsearch  
2015-07-22 MPU Document configuration and user experience of Fail2ban Added to MPU to-do list
2015-07-22 MPU Produce documentation on how to locate packages under the new setup How to find packages
2015-0610 MPU What does sync mean on a KVM client? On MPU list
2015-0610 MPU Consider descriptive KVM pool names On MPU list
10/12/14 MPU Investigate how local facilities for syslogd should be allocated On MPU list
22/10/14 MPU Compile notes on dumpdeps How to use dumpdeps
10/09/14 MPU MPU should consider adding MSI fix for HP DL180 bonding problems to appropriate server headers Done
09/04/14 MPU Move the BMCs of all KB servers to the new KB server management subnet Stephen has moved BMCs, Ian confirmed
28/08/13 MPU Create list of steps to take to 'un-manage' a DICE machine On mp-unit's ToDo list
08/05/13 MPU consider additional wake features such as a command to wake individual lab machines On their list
14/11/12 MPU Add "wake-ACL review" to to-do list. ACL modified
28/03/12 MPU Investigate removing pam_kx509 module by default from PAM stack (from deferred) Now on to-do list
28/03/12 MPU Produce apacheconf headerfile to switch on connection limits Done; see LCFG:core/include/dice/options/apacheconf-iplimit.h and LCFG:core/include/lcfg/options/apacheconf-iplimit.h
8/06/11 MPU Flesh out framework for handling firmware updates  
31/08/11 MPU Check upgrade to SL5.6 complete
17/08/11 MPU Grub and pxe passwords  
8/06/11 MPU make apacheconf-2 the default for SL6  
23/02/11 MPU Set up perpetual mirroring of latest CD image (LCFG disaster recovery)  
23/02/11 MPU check whether condor needs a kernel patch Not necessary in view of above
27/01/10 MPU Consider periodic file reconfiguration (om file configure) issues On mp-unit's list
22/02/17 MPU / Inf hosts / DNS ordering Now on MPU list
2015-09-09 MPU / RAT Review and update SL7 release notes  
11/03/15 MPU, RAT Consider likely FibreChannel requirements Services unlikely to expand. 8Gbps would be good to have. 6 per fabric.
27/07/16 MPU/Alastair Document SL7 KVM migration issues? Done, in SimpleKVMDocs and SimpleKVMHost
08/03/2017 MPU/Inf Look into time sync problem on VMs - possibly create script to report VMs with wrong time script done - no general problem noticed
2015-10-14 MPU/Inf Wire headers &c. - Discuss points raised in section 4 of Inf Unit report - e.g. eth0 or not?; &c. MPU believe sufficiently discussed. eth0 point will have to wait on NG network component
2015-10-14 MPU/Inf lcfg-dns - Discuss points raised in section 6 of Inf Unit report  
2015-09-23 MPU/Inf named on SL7 See the inf-unit report MPU/Inf to meet to discuss.
22/10/14 MPU/Inf Investigate AT network problem - why is data being lost/corrupted see RT 68813
2015-06-10 MPU/RAT Check current rootmail situation Neil's figures are here.
01/04/20 Neil Update VNC documents  
25/03/2020 Neil Start computing.help page on debugging remote working Initial page at http://computing.help.inf.ed.ac.uk/network-tests
fail2ban meeting Neil Look at logging all Apache logs centrally Fixed problem of logging to multiple places, but better done by project 540
11/09/19 Neil Ask Dave H. to look into sound deadening screens for AT-7.07 Ticket in forum issues RT 97845
27/02/2019 Neil Investigate how O365-using Schools handle root mail RootMailOnO365Summary
10/01/18 Neil Create project to cover move of http: sites to https: Done Project 454
13/09/17 Neil Coordinate our thoughts on the "open area" furniture  
08/03/2017 Neil Speak to Jon about request for web address - bayes.ed.ac.uk Done - DNS delegated to us
28/09/2016 Neil Check and mail out details of disk install issue Done - post to cos - basically how do you make sure which disk a machine installs to
11/11/15 Neil Agreed suitable mail log retention period with RAT and update wiki retention page with same Done, and updated http://computing.help.inf.ed.ac.uk/DP-statement-managed-systems
13/05/15 Neil Point "Contact Us" link at non-DICE support form Done
26/11/14 Neil Take "web pages with DP issues" to a TDM Did this as an AOCB on the 26/11/14
26/03/14 Neil Try to unbung mail Probably due to our relay being down. How robustly is that handled? Perhaps the action should be to make sure that improving the resilience of mail.inf is on the unit todo list. Added to ServicesUnitCurrentTasks
09/04/14 Neil check whether accessing mailman via the https interface allows viewing of other unit's mailing lists I have a working patch that adds a entitlements field to the list, people with that entitlement can admin the list. Configuration audit trails? - No, there are none.
12/03/14 Neil Talk to alastair about local homedirs Spoke to Alastair (on bus) he confirms we do still support local home dirs for users
26/06/13 Neil AFSDB to AT Done. afsdb0 on afsdbvmkb, afsdb1 on afsdbvmat, afsdb2 on fenrir
30/10/13 Neil Look into scraping of penetration testing results Had a look, not really feasible, too much javascript. Dumping of CSV file may be enough to view issues.
25/09/13 Neil Check Wordpress installations on homepages and groups.inf Only 1 remaining active WP on homepages and that's now been updated. Groups, none found (other than my test one)
12/06/13 Neil Ensure our drupal requirements are known Miles has now left and didn't send list. Neil now has a meeting with the web people on Monday 16th
14/11/12 Neil Arrange ifevo work ASAP.  
26/09/12 Neil Mail-flood discussion: think about where to go next with root-mail added to Services to-do list.  
28/03/12 Neil Add additional column to cookie list identifying responsibility for that cookie Updated ServicesUnitCookiesAudit with column
26/10/11 Neil Investigate/report pcounter charging oddities  
27/07/11 Neil add support for .sslaccess files to the School's webservers Done
22/06/11 Neil Produce documentation on how to Cosign protect pages Done
13/10/10 Neil Circulate list of servers which allow non-CO logins For amd removal
22/09/10 Neil Speak to Alison about Staffmail forwarding  
8/9/10 Neil Make macros for putting criticallity information into netgroups more generally available  
24/06/09 Neil Produce new guidelines on how to work around the issues with AFS and ssh public keys Updated "top 10"
25/02/15 Neil/Alastair consult records management about archived home directory RM responded.
25/04/12 Neil/Alison Investigate mining data previously mounted on diglett half the data's been rsync to cameleopard/rmirror21 and the other half to lammasu/rmirror18
2015-06-10 Neil/Chris Update not-a-service documentation Updated
23/11/11 Neil/Ian Investigate the loss of the jabber service Done. The suggested test (to remove the AFS cwd under which jabberd had been started) didn't cause any problem with the Jabber service, so we are currently none the wiser.
Question: to be clear about what happened two weeks ago: was it the case that the Jabber service completely broke? Can we make a detailed note of the actual symptoms and effects that were seen at the time?
14/11/12 Neil/Stephen Discuss AFS/package dependencies.  
06/04/11 Neil/Stephen Look at RAID1 lockups Behaviour not unexpected
08/10/14 Neil/Tim Consider alternative cgi server We did have a chat. Neil will check the current write access to the cgi directory, and Tim will put moving the affected exam mark pages to a dedicated VM
26/09/12 Neil/Tim Raise posting to DB-generated teaching lists with ITO  
12/07/17 Neil/Toby Have chat about ciphers ApacheCipherChat20180222
26/10/11 Neil/Toby Prometheus initial AFS volume creation  
11/12/13 neilb Investigate management of Wordpress See here
10/09/08 neilb Take "recycling of old machines" to ITPF There's a talk coming up...
13/01/2106 RAT Review rack use in IF, AT, KB  
26/11/14 (recurrent) RAT Review nagios use. The objective is to minimize the numbers of machines/services which are in permanently "error'ed-and-ack'ed" states. Refer to the current Nagios 'comments' page. Reminder: CEG has decreed that systems should be fixed promptly or removed from monitoring.
11/05/2016 RAT All Support team members should have access to Visitor tab in Theon Done.
13/05/15 RAT Update SSL certificates beast.inf, issrt.inf, rt4.inf - completed
09/12/15 RAT Check dufferin's root mail Done
27/01/2016 RAT Lecturer/TA-created VirtualBox images? Moved to RAT to-do list
12/11/14 RAT Look into producing seminar specific email lists for staff/students/etc tracked in TRAC
08/05/13 RAT Update computing.help to emphasise that students needing significant processing power for their projects should preferably use ECDF  
28/03/12 RAT Investigate issue of matlab status web page being tracked by RAT as part of flexlm component ap
23/11/11 RAT Tidy junk out of AT server room racks  
18/05/11 RAT check for package issues with making 64-bit default  
06/04/11 RAT Adjust pathfix configuration part of sl6 re-packaging
09/02/11 RAT split ipfilter.h Over to RAT for testing
23/02/11 RAT Investigate condor usage No-one appears to be using it so has been turned off
10/11/10 RAT Annual check of all open firewall holes for self-managed machines Done
10/11/10 RAT To look into Rootmail appearing to be polluted by the webots component  
28/07/10 RAT Units to set sysinfo.criticality if default of low is not appropriate  
13/01/10 RAT Ensure shutdown of beowulf cluster is more manageable Tested but needs someone else to test
8/9/10 RAT, Inf, US units Meet to discuss account issues Inf to convene
2015-06-24 RAT, US Review rack use in IF, AT, KB Forum 'non-fibre' racks (i.e. racks 6-14 - see the map) are particularly full. baphomet, ratte, stakhanov to be removed 13/11, roseval (now removed), scargill to be virtualised
13/01/2016 RAT/Inf Subnets for clusters Tracked at CEG, using wire 'O' for now
27/01/2016 RAT/MPU Student lab headers  
27/06/12 RAT/Toby Discuss grace periods with new DOT discussed at meeting held on 10/7 - can wait until lifecycle code
24/11/10 RAT? Create project to review use/status of bugzilla  
11/11/15 Roger Document for users the ability to scan to USB EnablingUSBonMFD (now user doc, linked from exam machine guide)
2015-07-22 Roger Take issue of package names to LCFG deployers meeting  
27/04/11 Roger Provide Stephen with apacheconf/cosign information  
23/06/10 Roger Take "user documentation" discussion points back to the project group done
15/01/14 Ross/Jennifer Pandemic actions Names now 'upper-cased'
22/01/2020 Services Investigate wiki speed Additional virtual hardware seems to have helped a bit
12/02/2020 Services Staffmail removal implications  
08/10/14 Services Trawl for pages restricted by ip address rather than via cosign No active ones in /teaching, leaving about 6 active ones elsewhere
fail2ban meeting Services Add fail2ban to git and gerrit  
08/10/14 Services Trawl for pages restricted by ip address rather than via cosign Neil identified 6 to be fixed - I've fixed the ones I'm going to
27/04/2016 Services Schedule WordPress upgrades on going - going to SL7 them at the same time
11/05/2016 Services Cloud printing - speak to other Schools about their experiences with Cloud printing  
09/12/15 Services Can users go negative with print credit? Yes, due to discrepancy between what software thinks the page count will be, and what actually gets printed by the printer. Workaround in progress. Being tracked in services todo list.
11/11/15 Services Make sure arrangements for visitor printing are sufficiently widely promulgated  
2015-09-09 Services Update computing.help re quotas Done
2015-07-22 Services Investigate measures to reduce blog comment spam http://computing.help.inf.ed.ac.uk/blog, also added a link to this page to new blog email, and first post page.
27/05/15 Services Make Yesterday point to readonly volumes It's on the list
27/05/15 Services Review Yesterday documentation Done - http://computing.help.inf.ed.ac.uk/yesterday
22/10/14 Services Investigate AFS aspect of AT network problem see RT 69292 - dropped due to inability to reproduce problem
27/11/13 Services From technical talk on mobile devices, add 'make websites more accessible' to Unit list or create project Done
27/11/13 Services From technical talk on mobile devices, add printing to Unit list or create project Done
27/02/13 Services Evo firmware upgrades icsa-evo done, leaving only ifevo2 to be replaced by ifevo4
14/03/12 Services Consider whether data location is optimum Moved to Services Unit task list
23/11/11 Services Change username/password for accessing disk arrays Done and details circulated to COs
14/09/11 Services Consider how to restrict printing in Forum (i.e. lab wires can't access Forum printers) .58 and .59 wires restricted
14/09/11 Services Enable pcounter on student servers (e.g. student.compute, student.login)  
13/07/11 Services Ensure pages under old groups documentation tree are unavailable Content still in CVS
09/03/11 Services Prevent onward logins from SVN server Add ssh-daemon (re)configuration to unit ToDo list
09/03/11 Services Move afsdb1 to different hardware if necessary Not necessary
14/04/10 services follow up on 24/02/10 AFS discussion: wrapper for long-running jobs, scripts server to be included in AFS enhancements project
14/12/16 Services Unit Arrange for removal of tape library from KB server room Removed!
28/09/2016 Services Unit Investigate options for allowing dynamic quotas to cope with large downloads DynamicAFSQuotasIssue1 - Raised at CEG. UG3/4/5 and MSc will get full quota from the start
08/06/2016 Services Unit Query users about NAS box purchases Several interesting points have emerged, it would appear that it's not all about cost.
08/06/2016 Services Unit Computing.help page for small NAS boxes http://computing.help.inf.ed.ac.uk/some-advice-buying-nas
12/11/14 Services Unit Change Tape Library power arrangements  
30/07/14 Services Unit Move final AFS DB server off of KDC Added to Services Unit todo list
30/10/13 Services Unit Add investigation of home directory structure to to-do list Added to ServicesUnitCurrentTasks
28/03/12 Services Unit Talk to IS about MDP print queues UniDesk I120409-0390 - I'd say done now
26/08/09 Services Unit Removal of AMD from servers RemovingAMDFromServers see action on ALL to act on Neil's list
26/08/09 Services Unit Removal of AMD from servers RemovingAMDFromServers see action on ALL to act on Neil's list
09/07/14 Services-unit Add bonding issues to BondingProblems MPU to investigate. See also this week's inf-unit report
26/10/11 services-unit Liaise with George/Malc re KB evo work RT:55508
27/04/11 services-unit Add ext2/3/4 to to-look-at list  
06/04/11 services-unit SAN partition monitoring Use /proc
12/08/09 Simon produce short note summarising the School's position on the use of EASE passwords with services and comment on use of AD subsumed by central doc
24/06/09 Simon Chase up production of documentation on how to install jabber clients on supported OSs Done
14/01/09 squinney Email a server example for auto reboot Done
22/10/08 squinney Arrange Christmas lunch Done
10/09/08 squinney Create install CD for Optiplex 755 Done
23/08/17 Stephen Enable apacheconf log compression fail2ban is on the MPU ToDo list
24/05/17 Stephen Move apacheconf logrotate "sharedscripts" to lcfg layer  
27/04/2016 Stephen Open area UG access rights Tighten up - done - uses new @login/forumpublic/console role
2015-06-10 Stephen Circulate dmesg proposal On the MPU list
23/04/14 Stephen Discuss pandemic "security" topic with interested parties  
09/07/14 Stephen Survey partition use and disc sizes and revise and expand EL7PartitionLayout  
28/09/11 Stephen Produce document on sanitising of local account UIDs and GIDs lcfg-level header now exists for uids/gids <= 700 - see LCFG wiki
09/11/11 Stephen Repeat AFS talks Scheduled for February
17/08/11 Stephen Look at reboot/halt questions  
27/04/11 Stephen Link component-version script  
26/01/11 Stephen Investigate creating a header file for self-managed machines with dynamic IP addresses On MPU list
12/01/11 Stephen produce worked svn examples, in particular branching  
13/10/10 Stephen Circulate lists of what's in the "devel" bucket  
10/06/09 Stephen send out mail detailing ancient entries in live_testing_defaults.rpms Done
23/08/17 Stephen / Graham /home and exam lockdown  
09/10/13 Stephen / Graham weka on NX weka removed from machines
fail2ban meeting Stephen/George Define range of local network addresses Some added to live/subnets.h and live/subnets6.h, but likely still incomplete
25/01/12 Stephen/Graham Discuss pam_console and exam machines DONE
14/11/12 Stephen/Neil/Toby perl-AFS package management.  
08/10/14 Support Clarify COs entitlement to Dreamspark agreement Entitlement confirmed but will write up - see DreamSpark
08/10/14 Support contact Records Management regarding the reconciliation of accounts. Discuss at meeting on 2nd December
23/11/11 Support All server room desktops should have local home directories Done, there will be a header for server room desktops shortly
13/08/08 support Chase down "temporary dhcp" usage Done; RT:37705
26/06/19 Tim Progress on remote resits  
25/02/15 Tim investigate why transmit and receive checksum unloading was disabled in the DICE virtualbox host configuration (It has now been enabled) Tim will investigate only if this becomes an issue in the future.
25/06/14 Tim inSpace robotarium Investigate concerns, then refer to CEG as appropriate.
26/09/12 Tim Check with ITO re updates to non-standard students' records Tim confirmed via email after meeting that action is complete.
17/08/11 Tim Post to sys-announce re database changes  
8/06/11 Tim Check with Stuart Anderson about teaching packages and 64-bit Done, 64bit will be the default in labs
09/03/11 Tim Take issue of Matlab usage to Ian Stark No change for moment but monitor situation
14/04/10 Tim Inform Stuart and/or Teaching Committee of news service news  
14/10/09 Tim Confirm what data on maelcum needs to be kept  
14/10/09 Tim Inform support about the 'round-robin' changes for ssh'ing to lab machines  
12/08/09 Tim Look into politics of managing hamburg Done
11/03/09 Tim and Alison Clarify exam procedures with the ITO Action adopted by CEG
09/12/15 Tim/Alison On-line exams and FRB closure Confident we now have sufficient skills. Phone contact with RAT is sufficient if required.
26/01/11 Tim/Alison Come up with a solution for the issue of student helpers having the "staff" role Staff role removed from student helpers
10/06/09 Tim/Alison talk to CSTR about BP servers Ongoing
26/08/09 Tim/Alison Fix inventory problems with HP desktops progress made but still problems with MDP machines - now fixed
26/08/09 Tim/Stephen Create separate LCFG branch for on-line exam machines completed and documentation now added to lab exam procedures
13/12/17 Toby Consider consequences of making "no-grace" the default some thoughts - PrometheusNoGraceDefault, move to CEG
10/05/17 Toby make list of existing functional accounts available. Done
23/08/17 Toby New student account emails Then a month allowed by password portal to fix up things forgotten
08/03/2017 Toby Create project for 'sorting' roles Done
11/11/15 Toby Write a roles-management document for discussion TheProblemWithRoles
25/02/15 Toby Document how to query roles on the prometheus server Done - started RolesFAQ
28/01/15 toby Investigate handling of kerberos credentials during SL7 install It looks like kdcregister doesn't store any credentials on disk, so this isn't an issue
12/11/14 Toby look into intermediate certificates Not possible at the moment but perhaps Summer 15.
TDM 21/05/14 Toby iOS Authentication Investigations complete. Written-up
23/04/14 Toby Look for unusual access patterns in cosign logs Toby explained that this had been quite difficult to do but he couldn't see anything obviously unusual.
27/11/13 Toby To action proposals from meeting to discuss whether revealing 'people' LDAP info. to EDLAN is OK. There are parts of ou=People which shouldn't be revealed outside Informatics. 11th Feb meeting decided that ou=People branch (but nothing else) of ldap tree may be accessed anonymously from EDLAN. Role information must be removed from people objects first. Toby to set up but will be tracked elsewhere.
27/11/13 Toby Experiment with test LDAP server/virtual DICE from home using EDLAN VPN Done - email sent to MPU 29/11
13/03/13 Toby Raise 32-bit Windows 7 issues on Heimdal list bugreport submitted. For details see minutes from 10/4/13
22/08/12 Toby Develop Alastair's Dice relationship diagram Merge updates and recirculate. latest here
28/11/12 Toby Review JANET certs re new charging model  
30/05/12 Toby Automatic running of ldapBuildAmdMaps Should be replaced by a Prometheus conduit or similar
25/01/12 Toby Add "cookie" text to Cosign and iFriend Done but suggestion to make text more obvious on weblogin. Now more obvious (it's got a box around it!)
23/11/11 Toby Email Graeme Wood about changes to EASE authentication Graeme has replied sayaing that no major changes are imminent
17/08/11 Toby Circulate notes on "roles" meeting Notes
22/06/11 Toby Investigate resurrection web interface to jabber  
24/06/09 Toby Add to the AFS FAQ details on how to use kerberised ssh on supported OSs Done
24/06/09 Toby Produce a web page detailing how to manually add signed server certificates when reinstalling services which use these certificates Done
10/06/09 Toby mail out about upcoming Cosign v3 upgrade Done
25/03/09 Toby Poll ITPF re. EASE principals IS are consulting
12/11/08 Toby Advance warning of loss of staffmail accounts, possibly via IDMS feed? Done
12/11/08 Toby Look at LDAP logging Done
10/09/08 toby Email timetable for installing new Informatics root certificate Done
27/04/2016 Toby / Stephen / Others Discuss fail2ban configurations Meeting scheduled - 18/08 10am, IF-1.15. All welcome. Fail2BanMeeting20160818
12/09/12 Toby/Alastair IS's information provision  
30/10/13 Toby/Graham Set up scripts area in CO group space /afs/inf.ed.ac.uk/group/cos/utils/
09/11/11 Toby/Graham investigate the LCFG client hang when switching to and from exam lockdown mode.
25/10/17 Toby/Jennifer Document procedure for creating roles/entitlements to grant access to individual machines. Done
28/05/14 Toby/Neil Making changes for ResetTheNet Some final changes after apache upgrade transferred to Services Unit to-do list
30/05/12 Toby/Neil New certificates Done 12/6/2012
13/05/09 Toby/Simon Think about afs principal (cf. chatroom) principal will be deleted
26/10/16 US Write a page on exFAT options  
26/10/16 US Notify cos of exam dates once known US report
27/04/2016 US Server room desktops Done
09/12/2015 US bioboy disks wiped, boxed and put in B.03
09/12/15 US Review SL7 tickets, and pass to MPU/RAT as required Upgrades have been restarted
27/11/13 US From technical talk on mobile devices, add AV to Unit list or create project Added to Unit list
09/01/13 US Ask a number of taught students whether they received email about Office365. All that were asked had received notification
14/03/12 US Produce folder of essential computing staff contact information Completed document in top drawer of 2.09
14/03/12 US Talk to research groups about server location Being done as part of SL6 upgrade
06/04/11 US Link systems blog from main "systems" page  
09/03/11 US Check out server rack rails in B.03 Done, though not labelled
23/02/11 US reboot remaining servers ASAP Just a few to do (btw, no need to reboot just for glibc changes)
09/02/11 US install VM student lab development machine Machine should be given more meaningful name
10/11/10 US Installing f13 on student.compute and student.login completed 23/2/2011
27/01/10 US Reconfigure ex-mars machines with root partition on large disk Done
08/11/17 US Unit Remove unrequired Tech access Done: the capability rfe/lcfg/write has been removed from the role techs. (See item 2 of https://www.dice.inf.ed.ac.uk/units/infrastructure/Projects/399-Second-factor/restricted/short-term-fixes-made.html)
16/05/12 US Unit Publish a list of public servers On new website
08/02/12 US Unit Investigate issues with non-authenticated DICE support form Neil has changed cgi to use cosign (where available) to populate username field.
23/03/11 US Unit Create web page for future interruptions to services Done, and linked from systems page
8/9/10 US Unit Print off criticality info once complete by next meeting
22/10/08 US unit Download SmartBoard drivers and make available to members of School Done
13/12/17 US-Unit Check self-managed status of CDT machines Checked. 7 DICE, 10 Linux. No-one is dual booting.
26/06/13 US-unit Server-room desktops to be "full CO"  
13/03/13 us-unit Deal with old forumtracker so VMware can be retired  
27/03/13 US/George Let Ian know which sysadmin tools you would like to see on servers The discussion can be seen at https://wiki.inf.ed.ac.uk/DICE/ExtraSysadminSwOnServers. This part is considered complete. US to document how to access the web interfaces of printers using tunnelling/authenticated proxy - see PrinterWebInterface. George to consider wireshark
26/09/12 US/MPU How to spot off-the-air machines sooner? On the MPU to-do list
24/09/14 User Support Investigate information given to new users about passwords Confirmed that text on pp.inf and in email sent to users needs to be changed. Confirmed that IS do not send anything that might cause confusion
09/04/14 User Support scan self managed machines with tool provided by Graham Done
28/04/10 User Support Mail users about introduction of fail2ban  
23/04/14 User-support Investigate twitter for system status reporting see infalerts still a little tweaking and documenting to do - See Ross for account password
31/08/11 User-support AT-printing notices and FAQ changes  
31/08/11 Yan Review past RT tickets from Unidesk perspective  

-- CraigStrachan - 28 Aug 2009

Topic revision: r760 - 02 Jul 2020 - 09:24:46 - CraigStrachan
 
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback
This Wiki uses Cookies